plz help me with this question although i have tried many times (in fourteen years time mother will be twice as old as her son, four years ago the sum of their ages was thirty. find the age of the mother when the sun was born?)
Answer by Cromlix(4381)

Hi there,
in fourteen years time mother
will be twice as old as her son.
M + 14 = 2(S + 14).......(1)
Four years ago the sum of their
ages was thirty.
M - 4 + S - 4 = 30.......(2)
Consider Equation(1)
M + 14 = 2(S + 14).......(1)
M + 14 = 2S + 28
M = 2S + 28 - 14
M = 2S + 14
Substitute (2S + 14) for M in Equation(2)
M - 4 + S - 4 = 30.......(2)
(2S + 14) - 4 + S - 4 = 30
2S + 14 - 4 + S - 4 = 30
Collect like terms:
2S + S = 30 - 14 + 4 + 4
3S = 24
S = 8
Using Equation(1)
M + 14 = 2(S + 14).......(1)
M + 14 = 2(8 + 14)
M + 14 = 2(22)
M + 14 = 44
M = 44 - 14
M = 30.
At present time:
Son is 8 years old
Mother is 30 years old
When the boy was born the mother's
age was 30 - 8 = 22 years old.
Hope this helps :-)